# Comparing Pre-trained Human Language Models: Is it Better with Human Context as Groups, Individual Traits, or Both?

### v2

@@ Abstract
Incorporating human context into language models is the next frontier for human-centered natural language processing. Currently, two pre-training methods exist: group-wise attributes (e.g., over-45-year-olds) or individual traits. Group attributes are coarse — not all 45-year-olds write the same way — while modeling individual traits allows for a more personalized representation, but requires more complex modeling and data. So far, it is unclear which pre-training approach benefits what tasks. We compare pre-training models with human context via 1) group attributes, 2) individual users, and 3) a combined approach on 5 user- and document-level tasks. We find that pre-training with both group and individual features significantly improves the two user-level regression tasks like age estimation and personality assessment. Pre-training on individual users significantly improves the three document-level classification tasks like stance and topic detection. It even does well for downstream tasks without historical user data. Our results suggest both approaches have specific use cases, opening new avenues for human-centered language modeling.
@@ 1 Introduction
Language varies between people. To capture this notion, two strands of human-centered Natural Language Processing (NLP) emerged to model the people behind the language. The first approach focuses on the group context, building on the sociolinguistic notion of specific socio-demographic attributes influencing the language of a particular group. These socio-demographic attributes include age, gender (Volkova et al., 2013; Hovy, 2015), location (Kulkarni et al., 2016; Garimella et al., 2017), personality (Schwartz et al., 2013; Lynn et al., 2017), and more. The second approach focuses on building personalized language models (PLMs) that target individualistic contexts (King and Cook, 2020; Delasalles et al., 2019), and latent attributes inferred from an individual’s historical language (Matero et al., 2021; Soni et al., 2022) to better model the user.
@@
While these two strands have advanced human-centered NLP, we still do not understand their relative strengths, complementarity, and impact on different tasks (Soni et al., 2023): People are not defined by their group membership alone Orlikowski et al. (2023), but individual traits might not generalize. We compare the downstream performance of models pre-trained with the two approaches as well as their combination. We answer the following broader research questions:
(RQ1): How can we incorporate group and/or individual human context into pre-training?
@@
(RQ2): How does pre-training with different human contexts differ in terms of downstream performance for different tasks?
We use the monolingual socio-demographically adapted model from Hung et al. (2023) and the HaRT model from Soni et al. (2022) for group and individual human contexts, respectively. For a model trained with both individual and group human context, we use a multi-task learning PLM based on HaRT to create a model, GRIT, in two variants: GRITage is adapted to the authors’ age, and GRITope to their inferred personality trait (openness). We test all systems on five user- and document-level tasks.
@@
Note that because we focus on empirically comparing pre-training group and individual traits in PLMs, we cannot compare to large language models like GPT4, which do not support stratification to either attribute via pre-training. Recent studies have explored methodologies like user-adapters (Zhong et al., 2021), and user-centric prompting (Li et al., 2023) to include human context into models. In contrast, we focus specifically on comparing the impact of pre-training LMs with different human contexts.
@@
We find that PLM pre-trained on individuals and group features enhances 2 user-level regression tasks: age estimation and personality assessment from multiple user documents. These user-level tasks focus on individual people. Our findings, that they are best modeled as a combination of their group attributes and unique characteristics, conforms with the notion that a person is a mix of both. Document-level categorization tasks, like stance detection, are more personal than group based. A PLM pre-trained on individual human context alone improves 3 document-level classification tasks.
@@
By their very nature, models of this kind touch upon sensitive user information. For this reason, we take a responsible release strategy, making only the code for the comparisons publicly available and the exact splits of the TrustPilot and Stance datasets used. We build on top of the publicly available code for HaRT and Hung et al. (2023). We acquired the model and data in a secure manner from the authors of Soni et al. (2022) and TrustPilot data splits from the authors of Hung et al. (2023). For more information about the model and data, see Sections 3 and 4. For a discussion of the ethical implications of the model and data, see Ethical Considerations section.
@@
(1) We compare pre-training several language models with individual users, group socio-demographic features, and both traits.
(2) We evaluate the three specific pre-training strategies on five downstream tasks: two multi-document user-level regression (personality-openness evaluation and age estimation) and three single document-level classification tasks (stance detection, topic detection, and age category prediction).
(3) We find that the two user-level regression tasks perform better when pre-trained with both strategies (GRIT) and the three single document-level tasks perform better on fine-tuning a language model pre-trained with individual context alone (HaRT).
@@ 2 Integrating Human Context in PLMs
For our comparison, we use three systems representing the paradigms of pre-training with human context (Figure 1). We want to tease apart the contributions of 1) grouping people, 2) modeling individual users, and 3) modeling both group and individual human contexts. As noted earlier, we focus on recent approaches for pre-training language models with the additional human context.
@@
We build on Hung et al. (2023), a recent model to explore demographic adaptation in transformer-based PLMs. They use bidirectional auto-encoder-based PLMs to inject demographic knowledge in a multi-task learning setup where they also train masked language modeling (MLM) and classify the gender or age of an author. They use the multilingual reviews dataset with demographic labels from Trustpilot [^1] https://www.trustpilot.com/ Hovy (2015). They evaluate multiple text classification tasks, including demographic attribute classification, sentiment analysis, and topic detection. We use the US-English subset of the Trustpilot data for topic detection (TD) under the age categories and for age attribute classification (AC) (more details in section 4). We compare to the results from Hung et al. (2023) with monolingual BERT on a US-English dataset with out-of-domain demographic (age) specialization. Out-of-domain data is the Blogs authorship corpus Schler et al. (2006), and in-domain means Trustpilot corpus. To be consistent and fair in comparing with other human context pre-training paradigms, we choose the monolingual model and eliminate the confounds from domain specialization.
@@
Soni et al. (2022) introduced human language modeling (HuLM) in pre-trained LMs. I.e., pre-training a regular language modeling but including a dynamic individual human context vector derived from the authors’ historical texts. This vector captures the human states in which the text was generated. It also adds coherence to texts generated by the same author. They introduce a Human-aware Recurrent Transformer (HaRT), an autoregressive PLM for human language modeling. They evaluate the effect of individual human context on language modeling and multiple user-level and document-level downstream tasks. We use the results from HaRT on the user-level tasks, age estimation and personality (openness) assessment, and on a document-level task, stance detection, for our comparisons study.
@@
We train a PLM to integrate both individual and group human context in language modeling pre-training. We extend HaRT by predicting an additional human group attribute in a dynamic muti-task learning setup similar to Hung et al.. To facilitate comparison, we combine the pre-training inducing the individual human context through the author’s language and inject the group context by predicting a group attribute of the author.
We choose downstream tasks from both prior works (Soni et al., 2022; Hung et al., 2023) to compare our three language models pre-trained with different levels of human contexts: two user-level regression tasks and three single document-level classification task.
@@ 3.1 Pre-training with individual human context
Soni et al. (2022) introduced HaRT to incorporate individual human context into PLMs. They use a 12-layered autoregressive GPT-2 based architecture with a modified self-attention computation at layer 2. This modification to the query vector now includes the individual human context via a dynamic user-state vector.
@@ 3.2 Pre-training with group human context
Hung et al. (2023) explore socio-demographic adapted BERT models to inject group human context into PLMs. We use the names BERTDS and BERTage-MLM to denote their demographic (age) specialization using the multi-task learning setup and demographic adaptation with masked language modeling respectively.
@@
Hung et al. (2023) train for domain adaptation using the masked language modeling ($L_{mlm}$) loss and for classifying demographic category using the binary cross-entropy loss ($L_{dem}$). To account for the homoscedastic uncertainty Kendall et al. (2018) of losses, they adopt a dynamic MTL objective for training with group human context. Homoscedastic uncertainty is a task-dependent weighting to derive a multi-task loss function that can optimally learn the weights and balance the impact of multiple loss functions. This approach accounts for the different scales across multiple losses.
$\tilde{L_{t}}=\frac{1}{2\sigma_{mse}^{2}}L_{t}+\log\sigma_{t}$
@@ 3.3 Pre-training with both individual and group human context
GRoup and Individual HaRT is a LM pre-trained for the HuLM task and a (continuous) socio-demographic attribute prediction regression task in a multi-task learning setup. The PLM uses the user-state vectors to predict the socio-demographic attribute of the user.
@@
Because of compute limitations, we chunk a user’s language history into blocks and process them in a single forward pass. Each block of text from a user results in a user-state vector. We use the average of the user-state vectors from each non-padded block of texts from an author to compute their final user-state representation. This representation is layer-normed and linearly transformed before making a continuous-valued prediction for the specific attribute.
We pre-train one model for the continuous attribute age (GRITage) and one for the continuous attribute personality type openness (GRITope), respectively. The models train on a regression loss for the attribute prediction regression tasks using mean squared error ($L_{mse}$), and a classification loss for the HuLM task using cross-entropy loss ($L_{ce}$). Similar to Hung et al. (2023), we account for the homoscedastic uncertainty Kendall et al. (2018) of losses. We use the joint loss for a continuous and discrete output as derived in Kendall et al. (2018) and compute our joint objective as follows:
@@
where, $\sigma_{ce}$ and $\sigma_{mse}$ are the variances of the task-specific losses.
In practice, we use log variance for numerical stability and use the adjusted loss calculation as follows:
$\exp^{-\eta_{ce}}L_{ce}+\eta_{ce}+\frac{1}{2}(\exp^{-\eta_{mse}}L_{mse}+\eta_{mse})$
where $\eta_{x}=\log\sigma_{x}^{2}$. We let $\sigma_{ce}$ and $\sigma_{mse}$ be learnable parameters for the model. We do not halve the log-term of the cross-entropy loss since we found it to perform better with our multi-task learning experiments.
Pre-training data.
Training.
@@ 3.4 Fine-Tuning
We fine-tune GRIT and HaRT for downstream document-level tasks. Each downstream task has a separate fine-tuned model that is initialized with the respective model’s pre-trained parameters and trained using the respective downstream task labels and an appropriate loss function. We also use the historical language of a user where available for any of the downstream tasks. We use the last predicted token’s representation to predict the label in document-level classification tasks. We experimented with fine-tuning GRIT for user-level regression tasks in multiple ways, including 1) similar to HaRT, by using the averaged user-state vectors from GRIT, 2) same as previous but fine-tuning only the history module, attribute prediction module, and the downstream task head, 3) freezing all the parameters of GRIT and fine-tuning the human attribute prediction module alone. However, we found continued training as described in section 3.5 to perform best.
We used the Optuna framework Akiba et al. (2019) for hyperparameter search, closely following the experimental settings in Soni et al..
@@ 3.5 Transfer Learning
We experiment with continuing pre-training GRIT models for each group attribute. To this end, we pre-train GRITage capped to 4 training blocks and use this pre-trained model to continue MTL with the HuLM task and predict personality (openness). We do the same for GRITope and continued MTL with predicting age.

### v3

@@ Abstract
Pre-trained language models consider the context of neighboring words and documents but lack any author context of the human generating the text. However, language depends on the author’s states, traits, social, situational, and environmental attributes, collectively referred to as human context (Soni et al., 2024). Human-centered natural language processing requires incorporating human context into language models. Currently, two methods exist: pre-training with 1) group-wise attributes (e.g., over-45-year-olds) or 2) individual traits. Group attributes are simple but coarse — not all 45-year-olds write the same way — while individual traits allow for more personalized representations, but require more complex modeling and data. It is unclear which approach benefits what tasks. We compare pre-training models with human context via 1) group attributes, 2) individual users, and 3) a combined approach on five user- and document-level tasks. Our results show that there is no best approach, but that human-centered language modeling holds avenues for different methods.
@@ 1 Introduction
Language is a fundamental form of human expression that varies between people. Pre-trained Language Models (PLMs) account for the textual context of neighboring words and documents but lack the human context of the author “generating" the language. However, language is highly dependent on the human context Soni et al. (2024), i.e., an author’s changing states Fleeson (2001); Mehl and Pennebaker (2003), traits, social, situational, and environmental attributes. For example, a person’s language differs when hiking (situation/environment) versus when feeling dejected (state) over a breakup (situation). It is essential to model the additional human context to better understand human language with PLMs Soni et al. (2024). Two strands of human-centered Natural Language Processing (NLP) emerged to model the people behind the language. The first focuses on the group context, building on the sociolinguistic notion of specific socio-demographic attributes influencing the language of a particular group. These socio-demographic attributes include age, gender (Volkova et al., 2013; Hovy, 2015), location (Kulkarni et al., 2016; Garimella et al., 2017), personality (Schwartz et al., 2013; Lynn et al., 2017), and more. The second approach focuses on building personalized language models that target individualistic contexts (King and Cook, 2020; Delasalles et al., 2019), and latent attributes inferred from an individual’s historical language (Matero et al., 2021; Soni et al., 2022) to better model the user.
@@
While these two strands have advanced human-centered NLP, we still do not understand their relative strengths, complementarity, and impact on different tasks (Soni et al., 2024). People are not defined by their group memberships alone Orlikowski et al. (2023), and individual traits might not be generalizable enough across groups. Further, cross-cultural psychology research (Hofstede and Bond, 1984) notes the importance of both individualism and collectivism and Soni et al. suggest a flexible interplay of these concepts when including human context in PLMs. We might expect models informed by both group and individual features to perform better, but no data exists on this. In this work, we empirically evaluate these hypotheses and compare the effects of including different types of human context in PLMs (i.e., groups context (collectivism), individualistic aspects (individualism), and a combination of both) on specific tasks. We answer the following broader research questions:
(RQ1): How can we incorporate group and individual human context into pre-training?
@@
(RQ2): How does pre-training with different human contexts affect downstream performance for different tasks?
Recent works trained PLMs with socio-demographic group context (Hung et al., 2023), and individual context (Soni et al., 2022). However, no empirical studies compare the impacts of different types of human contexts included in language modeling. We compare the downstream tasks’ performance of models from these works with that of a new PLM trained with group and individual contexts. We test all systems on five downstream tasks from these works to ensure a variety of tasks and prediction properties at three levels: 1) user level, with historical language from authors (age estimation and personality assessment), 2) document-level, with historical language from some authors (stance detection), 3) document-level, without historical language from authors (topic detection and age category classification).
@@
Note that because we focus on empirically comparing pre-training with different types of human context, we cannot compare to the larger language models like GPT4, which are not yet pre-trained/trainable with human context but are envisioned to become large human language models in the future (Soni et al., 2024). Recent studies have explored methodologies like user-adapters (Zhong et al., 2021) and user-centric prompting (Li et al., 2023) to include human context into the larger language models. In contrast, we focus specifically on comparing the impact of pre-training LMs with different human contexts, as Soni et al. (2024) argue that pre-training will allow for modeling a richer human context by explicitly handling the multi-level structure of documents connected to people, as opposed to specific and limited benefits of user-centric prompting and fine-tuning (Salemi et al., 2023; Choi et al., 2023).
@@
PLMs pre-trained on individual and group features enhance performance on two user-level regression tasks that use multiple user documents: age estimation and personality assessment. In contrast, PLMs pre-trained on individual human context alone improve performance on document-level classification tasks like stance and topic detection. Our findings suggest user-level tasks focusing on individual people are best modeled as a mix of both group attributes and unique characteristics. However, document-level tasks that are more personal, like stance detection, are best addressed by modeling the individual context alone. Adding group attributes degrades performance.
@@
By their very nature, models of this kind touch upon sensitive user information. For this reason, we adopt a responsible release strategy, making only the code for the comparisons publicly available, along with the exact splits of the TrustPilot and Stance datasets used [^1] https://github.com/soni-n/HumanContextLanguageModels_Comparison. We build on top of the publicly available code from Soni et al. (2022) and Hung et al. (2023). We acquired the models and data from the authors of the respective works in a secure manner. For more information about the models and data, see Sections 4 and 5. For a discussion of the ethical implications of the models and data, see the Ethical Considerations section.
@@
(1) We empirically compare three pre-training strategies for language models with human context: individual traits, group socio-demographic features, and a combination of both.
(2) We evaluate each pre-training strategy on five downstream tasks: two multi-document user-level regression (personality-openness evaluation and age estimation) and three single document-level classification tasks (stance detection, topic detection, and age category classification).
(3) We find that the two user-level regression tasks perform better with PLMs pre-trained with individual and group human contexts. Conversely, the three single document-level tasks perform better with PLMs pre-trained with individual context alone. Further, pre-training with group and individual contexts reduces performance for document-level tasks.
2 Related Work
Socio-demographic and latent human factors.
Much work in human-centered NLP has focused on identifying and evaluating inclusion of human context in our models. Initial studies show benefits of grouping people by socio-demographic factors like age or gender (Volkova et al., 2013; Hovy, 2015) and geographical region (Bamman et al., 2014; Garimella et al., 2017) to capture the variation in language usage and meaning among different groups, and improving text classification tasks like sentiment analysis. Additionally, adapting to socio-demographic user factors (Lynn et al., 2017), social networks (Huang et al., 2014; Radfar et al., 2020), and social media attributes (Bamman and Smith, 2015) have been effective to enhance the performance in tasks like sarcasm detection, and toxic language detection. Some studies go beyond modeling explicit groups, to learn individual representations latently Jaech and Ostendorf (2018); Delasalles et al. (2019) or via historical language Matero et al. (2021).
Pre-traing with human context.
With respect to pre-trained LMs, recent studies have used adapter-based methodology (Li et al., 2021; Zhong et al., 2021) to include individual human contexts for downstream tasks. More recently, large language models have used user-centric prompting (Li et al., 2023) to include human context and evaluate on personalized and social tasks, resulting in mediocre performance (Salemi et al., 2023; Choi et al., 2023). However, few studies have explored including human context within the pre-training regime of LMs. Hung et al. (2023) generalize the task-specific EMPATH-BERT (Guda et al., 2021) to create a PLM injected with demographic group information using a dynamic multi-task learning setup. Additionally, Soni et al. (2022) pre-train a LM with individual human context derived from user’s historical language. Our study aims at comparing the impacts of pre-training LMs with individual, or group, or combined individual plus group human context.
@@ 3 Integrating Human Context in PLMs
For our comparison, we use three systems representing the three paradigms of pre-training with human context (Figure 1). We want to tease apart the contributions of different types of human context: 1) grouping people, 2) modeling individual users, and 3) modeling both group and individual human contexts. As noted earlier, we focus on recent approaches for pre-training language models with additional human context.
@@
We build on a model from Hung et al. (2023) that explores demographic adaptation in transformer-based PLMs. It is a bidirectional auto-encoder-based PLM injecting demographic knowledge in a multi-task learning setup where they train for masked language modeling (MLM) and classify the gender or age of an author. They use the Trustpilot dataset  [^2] https://www.trustpilot.com/ of multilingual reviews with demographic labels Hovy (2015), and evaluate on multiple text classification tasks, including demographic attribute classification, sentiment analysis, and topic detection. For our comparison study, we use the US-English subset of the Trustpilot data for two tasks: topic detection (TD) across two age categories, and age attribute classification (AC) (more details in section 5). Additionally, we use a monolingual BERT pre-trained with age specialization on the Blogs authorship corpus Schler et al. (2006). This choice allows us to eliminate a domain influence (i.e., Trustpilot reviews), given that the other PLMs under comparison lack this specialization.
@@
Soni et al. (2022) introduced human language modeling (HuLM) in PLMs, which is regular language modeling given an additional context of the individual generating the language. This additional context is a dynamic vector derived from the authors’ historical texts motivated by the idea of capturing the changing human states expressed in language. It also adds coherence to texts generated by the same author. They introduce Human-aware Recurrent Transformer (HaRT), an autoregressive PLM to evaluate the effect of individual human context on language modeling and multiple user-level and document-level downstream tasks. We build on HaRT and use two user-level tasks, age estimation and personality (openness) assessment, and on a document-level task, stance detection, for our comparisons study.
@@
We train a PLM to integrate both individual and group human context by introducing a multi-task learning setup into HaRT that incorporates group features. This approach facilitates training a PLM with both group and individual context. We evaluate the model on two multi-document user-level regression tasks: age estimation and personality assessment, and three single document-level classification tasks: stance detection, topic detection, and age group classification. Importantly, the only difference in this multi-task learning setup compared to HaRT is the inclusion of a demographic attribute prediction during pre-training, similar to how Hung et al. (2023) adapted traditional PLMs for group context.
@@ 4.1 Pre-training with individual human context
Soni et al. (2022) use a 12-layered autoregressive GPT-2 based architecture with a modified self-attention computation at layer 2. This modification to the query vector now includes the individual human context via a dynamic user-state vector.
@@ 4.2 Pre-training with group human context
Hung et al. (2023) explore socio-demographic adapted BERT models to inject group human context into PLMs. We use two models: one specialized for age (demographic attribute) under the multi-task learning setup, and the other adapted to the age corpora through standard masked language modeling. We denote these as BERTDS and BERTage-MLM, respectively.
@@
Hung et al. (2023) train for both domain adaptation using the masked language modeling ($L_{mlm}$) loss and for classifying demographic category using the binary cross-entropy loss ($L_{dem}$). Both losses must be combined to simultaneously learn multiple objectives. To account for the homoscedastic uncertainty Kendall et al. (2018) of both losses, they adopt a dynamic multi-task learning (MTL) objective for training with group human context. Homoscedastic uncertainty is a task-dependent weighting to derive a multi-task loss function that can optimally learn the weights and balance the impact of multiple loss functions and their different scales. The tasks are dynamically weighted using the variance of the task-specific loss ($\sigma_{t}^{2}$) over training instances of the task $t\in\{mlm,dem\}$:
$\tilde{L_{t}}=\frac{1}{2\sigma_{t}^{2}}L_{t}+\log\sigma_{t}$
@@ 4.3 Pre-training with both individual and group human context
We train HaRT under a multi-task learning setup for both the individual context — through the HuLM pre-training task (see Section 4.1) — and the group features — via a regression task to predict a (continuous) socio-demographic attribute of the author. We call the model as GRoup and Individual HaRT (GRIT). The model uses the user-state vectors (see Section 4.1) to predict the socio-demographic attribute of the author:
@@
We chunk a user’s language history into blocks and process them in a single forward pass. Each block of text from a user results in a user-state vector. We use the average of the user-state vectors from each non-padded block of texts from an author to compute their final user-state representation. This representation is layer-normed and linearly transformed before making a continuous-valued prediction for the specific attribute.
We pre-train one model for the continuous attribute age (GRITage) and one for the continuous attribute personality type openness (GRITope). The models train on a regression loss for the attribute prediction regression tasks using mean squared error loss ($L_{mse}$), and a classification loss for the HuLM task using cross-entropy loss ($L_{ce}$). We must combine both losses to jointly learn the two objectives and account for the homoscedastic uncertainty Kendall et al. (2018) of the losses. Since we combine a regression and a classification loss, we train the model to learn to balance the loss for a continuous and discrete output as derived in Kendall et al. (2018) and compute our joint objective as follows:
@@
where, $\sigma_{ce}^{2}$ and $\sigma_{mse}^{2}$ are the variances of the task-specific losses over the training instances of the respective tasks.
To add numerical stability, we adjust the loss calculation to use log of the variance:
$\exp^{-\eta_{ce}}L_{ce}+\frac{1}{2}(\exp^{-\eta_{mse}}L_{mse}+\eta_{ce}+\eta_{mse})$
where $\eta_{x}=\log\sigma_{x}^{2}$ for $x\in\{mse,ce\}$. We let $\sigma_{ce}$ and $\sigma_{mse}$ be learnable parameters for the model. In practice, we do not halve the $\eta_{ce}$ term in the above equation since we found it to perform better with our multi-task learning experiments.
Pre-training Data and Training.
We use the same Facebook posts dataset (Park et al., 2015) and training, validation, and test splits as those used by Soni et al. (2022). For both GRITage and GRITope, we use the demographic and personality scores, respectively, obtained from consenting Facebook users (Kosinski et al., 2013). This data is identical to that used by HaRT for the age estimation and personality assessment tasks. During training, we use a learning rate of 5e-5 in the multi-tasking training setup, employing the homoscedastic loss computation method described earlier. Following the experimental settings for HaRT, each training instance is capped to 4 blocks of 1024 tokens each. We use a train batch size of 1 per device and an evaluation batch size of 20 per device, trained over 2 GPUs for 8 epochs. Further details can be found in Appendix A.1.
@@ 4.4 Fine-Tuning
We utilize the results of fine-tuned BERTDS and BERTage-MLM from Hung et al. (2023), as well as fine-tuned HaRT models from Soni et al. (2022) where available. We fine-tune both GRIT models for all downstream tasks, and HaRT for 2 document-level tasks. Additionally, we use the Optuna framework Akiba et al. (2019) for hyperparameter search, closely following the experimental settings in Soni et al. (2022). Details can be found in Appendix A.2.
@@ 4.5 Transfer Learning
We experiment with fine-tuning GRITage in a multi-task learning setup for both the HuLM task and predicting personality (openness). Similarly, we fine-tune GRITope to predict age while also training for the HuLM task. We observe that this form of transfer learning yields the best performance for the user-level regression tasks (refer to Section 6.1).